# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### v3

@@ 1. Introduction
Let $R$ be a ring spectrum. A complex orientation of $R$ is an element $c\in R^{2}(\mathbb{CP}^{\infty})$ whose restriction along the map $\mathbb{C}\mathbb{P}^{1}\xhookrightarrow{}\mathbb{CP}^{\infty}$ corresponds to $1$ under the natural identification $R^{2}(\mathbb{C}\mathbb{P}^{1})\cong\pi_{0}(R)$. Geometrically, this can be regarded as a generalized Chern class: it is a degree $2$ $R$-valued characteristic class for complex line bundles, and the restriction condition calibrates it to behave well under tensor product (which is necessary for it to induce a formal group law). As chromatic homotopy theory has developed, however, it has become clear that it is not only a computational tool for understanding geometric objects, but also a sort of analogue of algebraic number theory for Brave New Algebra. This motif appears in Balmer’s interpetation of the Thick Subcategory Theorem ([1],[2]), for example, and Lurie’s ([5]) and Gregoric’s ([3]) work on the moduli stacks of oriented elliptic curves and formal groups (respectively), as well as more speculative work such as Salch’s “topological Langlands program” ([9],[10]). It would be highly desirable to have an interpretation of complex orientations purely in terms of this higher algebra, but one is presently lacking.
@@
The main idea of the theorem is that the map $\mathbb{C}\mathbb{P}^{1}\to\mathbb{CP}^{\infty}$ is a sort of transposed Hurewicz map, using the equivalences $\mathbb{C}\mathbb{P}^{1}\cong S^{2}$ and $\mathbb{CP}^{\infty}\simeq K(\mathbb{Z},2)$. Explicitly, this arises from the fact that $\mathbb{CP}^{\infty}$ is the free strict abelian group on the pointed space $S^{2}$ ([5]). This allows us to identify the map defining complex orientations with a certain map defined purely in terms of $\mathbb{E}_{2}$- and $\mathbb{E}_{1}$-structures on the relevant spaces.
@@
If a ring spectrum $R$ has a higher complex orientation, it is supported at chromatic heights 00 and $\infty$.
Essentially, $R$ has a $p$-complete part (for each $p$) and a rational part, and nothing else. This theorem is proven using Ravenel and Wilson’s computation of the $K(n)$-cohomology of Eilenberg-MacLane spaces ([8]).
@@
While this theorem is stated in terms of chromatic support, it is equivalent to describe it in terms of telescopic support. This is because, by a theorem of Hovey ([4]), $T(n)$- and $K(n)$-localization are equivalent for complex-oriented ring spectra.
In the context of Lurie’s theory of orientations (section 4.3 of [5]), these results help to explain the central role played by even-periodicity in chromatic homotopy theory. An $\mathbb{E}_{\infty}$-ring $R$ is complex-periodic if and only if there is an oriented formal group over $R$, i.e. if there exists a formal group $\widehat{G}$ over $R$ with a preorientation $e\in\pi_{0}\mathrm{Map}_{*}(S^{2},\Omega^{\infty}\widehat{G}(\tau_{\geq 0}R))$ whose Bott map is an equivalence. This works because $\mathbb{CP}^{\infty}$ is the free strict abelian group on $S^{2}$, so maps from $S^{2}$ into formal groups will factor through $\mathbb{CP}^{\infty}$. After linearization, provided that the Bott map is an equivalence, this factorization corresponds to the lifting in the definition of complex orientation. The “$1$-periodic” version of this is trivial because $S^{1}\simeq K(\mathbb{Z},1)$; the $2$-periodic version can be described in terms of $\mathbb{E}_{2}$-strictifications by Theorem A; and the $n$-periodic version for $n\geq 3$ (which would involve factoring through the map $S^{n}\to K(\mathbb{Z},n)$) will not occur in cases of chromatic interest by Theorem B. Thus the strictification process allowing us to describe the behavior of these rings in terms of formal groups only works in degree $2$.
@@ 1.1. Conventions
  • We write $\mathrm{Map}_{\mathbb{E}_{n}}(X,Y)$ to denote the mapping space in the category of (grouplike) $\mathbb{E}_{n}$-spaces. If we just write $\mathrm{Map}(X,Y)$ with no subscript, this denotes the mapping space in the category of spaces.
@@ 2. Strict Elements
We are therefore lead to ask: what is an “element” of a spectrum $X$? When $X$ is a space, we might reasonably use this to refer an element of $\pi_{0}X$. In the spectral case, however, this definition is problematic. Part of the problem is philosophical: this fails to take into account the structure of a spectrum beyond its infinite loop space. In particular, when $X$ is a connective spectrum identified with a grouplike $\mathbb{E}_{\infty}$-space, an element in this sense completely ignores the $\mathbb{E}_{\infty}$ structure. However, this notion of element also has practical issues, as described in the following motivating example (c.f. the first chapter of [5]).
@@
We have a problem here: $\mathbb{S}\{t^{\pm}\}$ is not flat over the sphere spectrum! Moreover, its homotopy groups are terribly unpleasant. For both of these reasons, its completion is not a formal group in the sense of Lurie, which prevents us from doing chromatic homotopy theory with it. Even though it is the most literal version of the units of $R$, it fails to behave like it should. This was the motivation behind Hopkins’s definition of the strict multiplicative group, $\mathbb{G}_{m}(R)\vcentcolon=\tau_{\geq 0}\mathrm{Map}_{\mathrm{Sp}}(H\mathbb{Z},gl_{1}(R))$. The functor $\mathbb{G}_{m}$ is corepresented by the $\mathbb{E}_{\infty}$-ring of Laurent polynomials in one variable, $\mathbb{S}[t^{\pm}]=\Sigma_{+}^{\infty}\mathbb{Z}$, which is flat over $\mathbb{S}$ and has nice homotopy groups $\pi_{*}\mathbb{S}[t^{\pm}])=(\pi_{*}\mathbb{S})[t^{\pm}]$.
@@
3. Higher Complex Orientations
We now place the theory of the previous section in a broader context. Observe that we have $\mathrm{Map}_{\mathrm{Sp}}(H\mathbb{Z},X)\simeq\mathrm{Map}_{\mathbb{E}_{\infty}}(\mathbb{Z},\Omega^{\infty}X)$ and $\pi_{0}X\simeq\mathrm{Map}_{\mathbb{E}_{1}}(\mathbb{Z},\Omega^{\infty}X)$. It is reasonable to consider notions of “partially strict” elements that interpolate between these:
@@ 3. Higher Complex Orientations
Knowing now that complex orientations are $\mathbb{E}_{2}$ lifts of $1$, a natural question to ask is what $\mathbb{E}_{n}$ lifts of $1$ look like for $n\geq 3$. These “higher complex orientations” turn out to be quite scarce, however. For the case $n=\infty$, an argument of Nardin and Peterson ([7]) gives us an obstruction.
@@
It is enough to show that the unit map of $K(n)\otimes R$ is zero, since a ring with $1=0$ is trivial. By Ravenel-Wilson’s computation of the Morava K-theory of Eilenberg-MacLane spaces ([8]), we find that $K(n)_{*}K(\mathbb{Z},3)$ is concentrated in even degree. It follows that the map $\mathbb{S}\to K(n)\otimes\Sigma^{\infty-3}K(\mathbb{Z},3)$ must be zero, and so, therefore, is the unit of $K(n)\otimes R$. ∎
@@
For the case $n=1$, $m=2$, this will evidently coincide with the cellular obstruction theory for extending from $\mathbb{C}\mathbb{P}^{1}\to\mathbb{CP}^{\infty}$. The case for greater $n$ and $m$ is less clear, however. By Dunn additivity ([6]), we have a canonical equivalence $\mathbb{E}_{n+1}(\mathrm{Spaces}_{*})\simeq\mathbb{E}_{1}(\mathbb{E}_{n}(\mathrm{Spaces}_{*}))$, so it should be possible to give an inductive description using $\mathbb{E}_{1}$ obstruction theory.

### v4

# ./refs.bib

@@ 1. Introduction
Let $R$ be a ring spectrum. A complex orientation of $R$ is an element $c\in R^{2}(\mathbb{CP}^{\infty})$ whose restriction along the map $\mathbb{C}\mathbb{P}^{1}\xhookrightarrow{}\mathbb{CP}^{\infty}$ corresponds to $1$ under the natural identification $R^{2}(\mathbb{C}\mathbb{P}^{1})\cong\pi_{0}(R)$. Geometrically, this can be regarded as a generalized Chern class: it is a degree $2$ $R$-valued characteristic class for complex line bundles, and the restriction condition calibrates it to behave well under tensor product (which is necessary for it to induce a formal group law).
As chromatic homotopy theory has developed, however, it has become clear that it is not only a computational tool for understanding geometric objects, but also a sort of analogue of algebraic number theory for Brave New Algebra. This motif appears in Balmer’s interpetation of the Thick Subcategory Theorem ([balmer2020],[barthel2020]), for example, and Lurie’s ([lurie2018]) and Gregoric’s ([gregoric2021]) work on the moduli stacks of oriented elliptic curves and formal groups (respectively), as well as more speculative work such as Salch’s “topological Langlands program” ([salch2023a],[salch2023]). It would be highly desirable to have an interpretation of complex orientations purely in terms of this higher algebra, but one is presently lacking.
@@
The main idea of the theorem is that the map $\mathbb{C}\mathbb{P}^{1}\to\mathbb{CP}^{\infty}$ is a sort of transposed Hurewicz map, using the equivalences $\mathbb{C}\mathbb{P}^{1}\cong S^{2}$ and $\mathbb{CP}^{\infty}\simeq K(\mathbb{Z},2)$. Explicitly, this arises from the fact that $\mathbb{CP}^{\infty}$ is the free strict abelian group on the pointed space $S^{2}$ ([lurie2018]). This allows us to identify the map defining complex orientations with a certain map defined purely in terms of $\mathbb{E}_{2}$- and $\mathbb{E}_{1}$-structures on the relevant spaces.
@@
If a ring spectrum $R$ has a higher complex orientation, then $L_{K(n)}R\simeq 0$ at all primes $p$ for each $0<n<\infty$.
That is to say, a ring spectrum with a higher complex orientation is supported at chromatic heights 00 and $\infty$. [^1] While this theorem is stated in terms of chromatic support, it is equivalent to describe it in terms of telescopic support. This is because, by a theorem of Hovey ([hovey1995]), $T(n)$- and $K(n)$-localization are equivalent for complex-oriented ring spectra. This theorem is proven using Ravenel and Wilson’s computation of the $K(n)$-cohomology of Eilenberg-MacLane spaces ([ravenel1980]).
@@
These results help to explain the central role played by even-periodicity in chromatic homotopy theory. An $\mathbb{E}_{\infty}$-ring $R$ is complex-periodic if and only if there is an oriented formal group over $R$, and this is ultimately the case because $\mathbb{CP}^{\infty}$ is the free strict abelian group on $S^{2}$. Using this relation shows that the Bott map of an orientation (see [lurie2018]) will factor through the map $S^{2}\to\mathbb{CP}^{\infty}$ to give the lifting in the definition of a complex orientation.
The “$1$-periodic” version of the story is trivial because $S^{1}\simeq K(\mathbb{Z},1)$; the $2$-periodic version can be described in terms of $\mathbb{E}_{2}$-strictifications by Theorem A; and the $n$-periodic version for $n\geq 3$ (which would involve factoring through the map $S^{n}\to K(\mathbb{Z},n)$) will not occur in cases of chromatic interest by Theorem B. Thus the strictification process allowing us to describe the behavior of these rings in terms of formal groups only works in degree $2$.
@@ 1.1. Conventions
  • We write $\mathrm{Map}_{\mathbb{E}_{n}}(X,Y)$ to denote the mapping space in the category of (grouplike) $\mathbb{E}_{n}$-spaces and $\mathrm{Map}_{*}(X,Y)$ to denote the mapping space in the category of pointed spaces.
@@ 2. Strict Elements
We are therefore lead to ask: what is an “element” of a spectrum $X$? When $X$ is a space, we might reasonably use this to refer an element of $\pi_{0}X$. In the spectral case, however, this definition is problematic. Part of the problem is philosophical: this fails to take into account the structure of a spectrum beyond its infinite loop space. In particular, when $X$ is a connective spectrum identified with a grouplike $\mathbb{E}_{\infty}$-space, an element in this sense completely ignores the $\mathbb{E}_{\infty}$ structure. However, this notion of element also has practical issues, as described in the following motivating example (c.f. the first chapter of [lurie2018]).
@@
We have a problem here: $\mathbb{S}\{t^{\pm}\}$ is not flat over the sphere spectrum! Moreover, its homotopy groups are terribly unpleasant. For both of these reasons, its completion is not a formal group in the sense of Lurie, which limits its usefulness in chromatic homotopy theory. Even though it is the most literal version of the units of $R$, it fails to behave like it should. This was the motivation behind Hopkins’s definition of the strict multiplicative group, $\mathbb{G}_{m}(R)\vcentcolon=\tau_{\geq 0}\mathrm{Map}_{\mathrm{Sp}}(H\mathbb{Z},gl_{1}(R))$. The functor $\mathbb{G}_{m}$ is corepresented by the $\mathbb{E}_{\infty}$-ring of Laurent polynomials in one variable, $\mathbb{S}[t^{\pm}]=\Sigma_{+}^{\infty}\mathbb{Z}$, which is flat over $\mathbb{S}$ and has nice homotopy groups $\pi_{*}\mathbb{S}[t^{\pm}])=(\pi_{*}\mathbb{S})[t^{\pm}]$.
@@
We can go further with this. Observe that we have $\mathrm{Map}_{\mathrm{Sp}}(H\mathbb{Z},X)\simeq\mathrm{Map}_{\mathbb{E}_{\infty}}(\mathbb{Z},\Omega^{\infty}X)$ and $\pi_{0}X\simeq\mathrm{Map}_{\mathbb{E}_{1}}(\mathbb{Z},\Omega^{\infty}X)$. It is reasonable to consider notions of “partially strict” elements that interpolate between these:
@@
3. Higher Complex Orientations
Knowing now that complex orientations are $\mathbb{E}_{2}$ lifts of $1$, a natural question to ask is what $\mathbb{E}_{n}$ lifts of $1$ look like for $n\geq 3$. These “higher complex orientations” turn out to be quite scarce, however. For the case $n=\infty$, an argument of Nardin and Peterson ([nardin2018]) gives us an obstruction.
@@ 3. Higher Complex Orientations
It is enough to show that the unit map of $K(n)\otimes R$ is zero, since a ring with $1=0$ is trivial. By Ravenel-Wilson’s computation of the Morava K-theory of Eilenberg-MacLane spaces (Theorem 12.1 of [ravenel1980]), we find that $K(n)_{*}K(\mathbb{Z},3)$ is concentrated in even degree. It follows that the map $\mathbb{S}\to K(n)\otimes\Sigma^{\infty-3}K(\mathbb{Z},3)$ must be zero, and so, therefore, is the unit of $K(n)\otimes R$. ∎
@@
For the case $n=1$, $m=2$, this will evidently coincide with the cellular obstruction theory for extending from $\mathbb{C}\mathbb{P}^{1}\to\mathbb{CP}^{\infty}$. The case for greater $n$ and $m$ is less clear, however. By Dunn additivity ([lurie2017]), we have a canonical equivalence $\mathbb{E}_{n+1}(\mathrm{Spaces}_{*})\simeq\mathbb{E}_{1}(\mathbb{E}_{n}(\mathrm{Spaces}_{*}))$, so it should be possible to give an inductive description using $\mathbb{E}_{1}$ obstruction theory.
[…]